# The Mechanistic Emergence of Symbol Grounding in Language Models

### v1

@@ 1 Introduction
With the rising of powerful autoregressive language models (LMs; OpenAI, 2024; Anthropic, 2024; Comanici et al., 2025, inter alia) and their VLM extensions, there is growing interest in identifying and interpreting their emergent capabilities. Recent work has shown preliminary correlational evidence that grounding may emerge in LLMs (Sabet et al., 2020; Shi et al., 2021; Wu et al., 2025b) and VLMs (Cao et al., 2025; Bousselham et al., 2024; Schnaus et al., 2025) trained at scale, even when solely optimized with the simple next-token prediction objective. However, the potential underlying mechanisms that lead to such an emergence are not well understood. To address this limitation, our work seeks to understand the emergence of symbol grounding in LMs, causally and mechanistically tracing how symbol grounding arises within the internal computations.
@@ 2 Related Work
2.2 Emergent capabilities and learning dynamics of LMs.
@@ 3.1 Dataset and Tokenization
Child-directed speech. The Child Language Data Exchange System (CHILDES; MacWhinney, 2000) provides transcripts enriched with environmental annotations.‡‡‡See the manual for data usage: https://talkbank.org/0info/manuals/CHAT.pdf We use the spoken utterances as the linguistic tokens (LAN) and the environmental descriptions as the environment tokens (ENV). The environmental context is drawn from three annotation types:
@@ 5.3 Causal Interventions of Attention Heads
5.4 Generalization to Visual Dialogue with Images

### v2

@@ 1 Introduction
With the rising of powerful autoregressive language models (LMs; OpenAI, 2024; Anthropic, 2024; Comanici et al., 2025, inter alia) and their VLM extensions, there is growing interest in identifying and interpreting their emergent capabilities. Recent work has shown preliminary correlational evidence that grounding may emerge in LMs (Sabet et al., 2020; Shi et al., 2021; Wu et al., 2025b) and VLMs (Cao et al., 2025; Bousselham et al., 2024; Schnaus et al., 2025) trained at scale, even when solely optimized with the simple next-token prediction objective. However, the potential underlying mechanisms that lead to such an emergence are not well understood. To address this limitation, our work seeks to understand the emergence of symbol grounding in LMs, causally and mechanistically tracing how symbol grounding arises within the internal computations.
@@ 2.1 Language Grounding
2.2 Emergent Capabilities and Learning Dynamics of LMs
@@ 3.1 Dataset and Tokenization
Child-directed speech. The Child Language Data Exchange System (CHILDES; MacWhinney, 2000) provides transcripts of speech enriched with environmental annotations.‡‡‡See the manual for data usage: https://talkbank.org/0info/manuals/CHAT.pdf We use the spoken utterances as the linguistic tokens (LAN) and the environmental descriptions as the environment tokens (ENV). The environmental context is drawn from three annotation types:
@@ 5.3 Causal Interventions of Attention Heads
5.4 Generalization to Visual Dialog with Images